\begin{thebibliography}{99}\small

\bibitem{A1999}
M. D. Acosta, \textit{Denseness of norm attaining mappings}, Rev. R. Acad. Cienc. Ser. A. Mat. \textbf{100} (2006), 9--30.

\bibitem{AAGM08}
M. D. Acosta, R. M. Aron, D. García, M. Maestre, \textit{The Bishop–Phelps–Bollobás theorem for operators}, J. Funct. Anal. \textbf{254} (2008), 2780--2799.

\bibitem{ACKLM15}
R. M. Aron, Y. S. Choi, S. K. Kim, H. J. Lee, M. Martín, \textit{The Bishop-Phelps-Bollobás version of Lindenstrauss properties A and B}, Trans. Amer. Math. Soc. \textbf{367} (2015), 6085--6101.

\bibitem{BP1961}
E. Bishop, R. R. Phelps, \textit{A proof that every Banach space is subreflexive}, Bull. Amer. Math. Soc. \textbf{67} (1961), 97--98.

\bibitem{B1970}
B. Bollobás, \textit{An extension to the theorem of Bishop and Phelps}, Bull. London Math. Soc. \textbf{2} (1970), 181--182.

\bibitem{CGKM14}
Y. S. Choi, D. García, S. K. Kim, M. Maestre, \textit{Some geometric properties of disk algebras}, J. Math. Anal. Appl. \textbf{409} (2014), 147--157.

\bibitem{CK12}
Y. S. Choi, S. K. Kim, \textit{The Bishop–Phelps–Bollobás property and lush spaces}, J. Math. Anal. Appl. \textbf{390} (2012), 549--555.

\bibitem{DGMM18}
S. Dantas, D. García, M. Maestre, M. Martín, \textit{The Bishop–Phelps–Bollobás property for compact operators}, Canad. J. Math. \textbf{70} (2018), no. 1, 53--73.

\bibitem{DGMO}
S. Dantas, D. García, M. Maestre, Ó. Roldan, \textit{The
Bishop-Phelps-Bollobás theorem: and overview.} Operators and Norm Inequalities and Related Topics.  (2022), Trends in Mathematics, 519--576.

\bibitem{KLM15}
S. K. Kim, H. J. Lee, M. Martín, \textit{The Bishop–Phelps–Bollobás theorem for operators from $\ell_1$ sums of Banach spaces}, J. Math. Anal. Appl. \textbf{428} (2015), 920--929.

\bibitem{KL22}
S. K. Kim, H. J. Lee, \textit{Denseness of norm attaining compact operators to some vector‑valued function spaces}, Banach J. Math. Anal. \textbf{16}, 69 (2022).

\bibitem{L1963}
J. Lindenstrauss, \textit{On operators which attain their norm}, Isr. J. Math. \textbf{1} (1963), 139--148.

\bibitem{M2016}
M. Martín, \textit{The version for compact operators of Lindenstrauss properties A and B}, Rev. R. Acad. Cienc. Ser. A. Mat. \textbf{110} (2016), 269--284.

\bibitem{S83}
W. Schachermayer, \textit{Norm attaining operators on some classical Banach spaces}, Pacific J. Math. \textbf{105} (1983), no. 2, 427--438.

\end{thebibliography}

\documentclass[12 pt]{article}
\usepackage{amsmath, amsthm, latexsym}
\usepackage[all]{xy}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{color, soul, mathabx}
\usepackage{graphicx}
\usepackage{authblk}
\usepackage[utf8]{inputenc}
\usepackage[hidelinks]{hyperref}
\usepackage{orcidlink}
\providecommand{\U}[1]{\protect\rule{.1in}{.1in}}

\newtheorem{theorem}{Theorem}[section]

\theoremstyle{definition}
\newtheorem{dfn}[theorem]{Definition}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
       
\newcommand{\conjC}{\mathbb{C}}
\newcommand{\conjR}{\mathbb{R}}
\newcommand{\conjN}{\mathbb{N}}
\DeclareMathOperator{\na}{NA }
\DeclareMathOperator{\interior}{int }

\title{Stability results of the Bishop-Phelps-Bollobás property and the generalized AHSP \thanks{The authors contributed equally to this work.}}

\author{\footnote{Corresponding author}\, $^1$Thiago Grando \orcidlink{0000-0002-7590-7358}, \footnote{Corresponding author}\, $^2$Elisa R. Santos \orcidlink{0000-0002-4575-3001}}

\affil{$^1$ \small Department of Mathematics, Midwestern Paraná State University, \\ Guarapuava, Paraná, Brazil.
\texttt{tgrando@unicentro.br}}

\affil{$^2$Institute of Mathematics and Statistics, Federal University of Uberl\^andia, \\ Uberlândia, Minas Gerais, Brazil.
\texttt{elisars@ufu.br}}

\date{ }

\begin{document}

\maketitle

\begin{abstract}
In this paper, we study the Bishop-Phelps-Bollobás property for operators (BPBp for short). To this end, we investigate the generalized approximate hyperplane series property (generalized AHSP for short) for a pair $(X,Y)$ of Banach spaces, which characterizes when $(\ell_1(X),Y)$ has the BPBp. We prove the following results. For a locally compact Hausdorff space $L$, if $(X, \mathcal{C}_0(L,Y))$ has the BPBp, then so does $(X,Y)$. Furthermore, if the pair $(X, Y)$ has the generalized AHSP and $\mathcal{L}(X,Z) = \mathcal{K}(X,Z)$, then the pair $(X, Z)$ also has the generalized AHSP, where $Z$ is one of the spaces $\mathcal{C}(K, Y)$, $\mathcal{C}_0(L, Y)$, or $\mathcal{C}_b(\Omega, Y)$, with $K$ a compact Hausdorff space and $\Omega$ a completely regular space.
\end{abstract}

\

\textsc{Keywords:} Bishop-Phelps-Bollobás property, generalized approximate hyperplane series property, Banach spaces of continuous functions, norm attaining operators.

\

\textsc{MSC 2020:} Primary 46B20; secondary 46B25, 46E40.

\section{Introduction} 

Let $X$ and $Y$ be real or complex Banach spaces. We denote by $B_X$ the closed unit ball, by $S_X$ the unit sphere, and by $X^*$ the topological dual of $X$. We write $\mathcal{L}(X,Y)$ for the Banach space of all bounded linear operators from $X$ into $Y$, endowed with the operator norm. An operator $T \in \mathcal{L}(X,Y)$ is said to be \textit{norm attaining} if there exists $x \in X$ with $\|x\| = 1$ such that $\|Tx\| = \|T\|$. We denote by $\na(X,Y)$ the set of all norm attaining operators from $X$ into $Y$. An operator $T \in \mathcal{L}(X,Y)$ is called \textit{compact} if the set $\overline{T(B_X)}$ is compact in $Y$. The space of all compact operators from $X$ into $Y$ is denoted by $\mathcal{K}(X,Y)$.

The classical Bishop–Phelps theorem \cite{BP1961} states that if $Y$ is one-dimensional, then $\na(X,Y)$ is norm dense in $X^*$. In the same article, the authors asked whether, in general, $\overline{\na(X,Y)} = \mathcal{L}(X,Y)$. Lindenstrauss \cite{L1963} showed that there exist Banach spaces $X$ and $Y$ for which $\na(X,Y)$ is not dense in $\mathcal{L}(X,Y)$. To obtain positive results regarding the question raised in \cite{BP1961}, he introduced properties $A$ and $B$ for Banach spaces. A Banach space $X$ is said to have \textit{property $A$} if $\na(X,Y)$ is dense in $\mathcal{L}(X,Y)$ for every Banach space $Y$. Similarly, a Banach space $Y$ has \textit{property $B$} if $\na(X,Y)$ is dense in $\mathcal{L}(X,Y)$ for every Banach space $X$. Among other results, Lindenstrauss proved that reflexive spaces have property $A$, as does $\ell_1$, and that $c_0$, $\ell_\infty$, and finite-dimensional polyhedral spaces are examples of Banach spaces with property $B$. We refer the reader to the survey papers \cite{A1999} and \cite{M2016} for further background on the topic of norm attaining operators.

Afterwards, Bollobás \cite{B1970} showed that if $X$ is a Banach space and $0 < \varepsilon < 1$, then for any $x \in B_X$ and $x^* \in S_{X^*}$ satisfying $\left|1 - x^*(x)\right| < \frac{\varepsilon^2}{4}$, there exist $y \in S_X$ and $y^* \in S_{X^*}$ such that $y^*(y) = 1$, $\|x - y\| < \varepsilon$, and $\|x^* - y^*\| < \varepsilon$. This result is known as the Bishop–Phelps–Bollobás theorem.

In 2008, Acosta et al. \cite{AAGM08} initiated the study of this theorem in the context of operators between Banach spaces $X$ and $Y$, and introduced the Bishop–Phelps–Bollobás property for operators.

\begin{dfn}\label{def-BPBP-operadores}
A pair of Banach spaces $(X, Y)$ is said to have the \textit{Bishop-Phelps-Bollobás property for operators} (BPBp for short) if, given $\varepsilon > 0$, there is $\eta(\varepsilon) > 0$ such that for every $T \in S_{\mathcal{L}(X,Y)}$ and $x_0 \in S_{X}$ satisfying
$$\Vert  T x_0 \Vert > 1- \eta(\varepsilon),$$
there exist an operator $S \in S_{{\mathcal{L}(X,Y)}}$ and a point $u_0 \in S_{X}$ such that
$$
\| Su_0 \| = 1, \quad \Vert u_0 - x_0 \Vert < \varepsilon \quad \text{ and } \quad \Vert S-T \Vert < \varepsilon.
$$
In this case, we say that the pair $(X, Y)$ has the BPBp with function $\varepsilon \mapsto \eta(\varepsilon)$.
\end{dfn}

We refer the paper \cite{DGMO} for developments regarding the Bishop-Phelps-Bollobás property for operators.

A property called approximate hyperplane series property (AHSP for short) was introduced by Acosta et al. \cite{AAGM08} to characterize Banach spaces $Y$ for which the pair $(\ell_1, Y)$ has the BPBp.

\begin{dfn}\label{ahsp}
A Banach space $Y$ is said to have the property \textit{AHSP} if, for every $\varepsilon > 0$, there exists $0 < \eta(\varepsilon) < \varepsilon$ such that for every sequence $(y_k) \subset S_Y$ and every convex series $\sum_{k=1}^\infty \alpha_k$ with
$$\left\| \sum_{k=1}^\infty \alpha_k y_k \right\| > 1 - \eta(\varepsilon),$$
there exist a subset $A \subset \conjN$, $y^* \in S_{Y^*}$, and a subset $\{z_k : k \in A\} \subset S_Y$ satisfying the following:
\begin{enumerate}
    \item[(1)] $\sum_{k \in A} \alpha_k > 1 - \varepsilon$,
    \item[(2)] $\|z_k - y_k\| < \varepsilon$ for all $k \in A$,
    \item[(3)] $y^*(z_k) = 1$ for every $k \in A$.
\end{enumerate}
\end{dfn}

This property holds in various Banach spaces, including finite-dimensional spaces, uniformly convex spaces, spaces with property~$\beta$, and lush spaces \cite{AAGM08,CK12}.

In 2015, Kim et al. \cite{KLM15} introduced the notion of the generalized approximate hyperplane series property (generalized AHSP for short) for a pair $(X,Y)$ of Banach spaces. This property characterizes when the pair $(\ell_1(X), Y)$ has the BPBp, where $\ell_1(X)$ denotes the Banach space of all sequences $(x_k)$ in $X$ such that the series $\sum_{k=1}^\infty\|x_k\|$ converges.

\begin{dfn}\label{gahsp}
A pair of Banach spaces $(X, Y)$ is said to have the \textit{generalized AHSP} if, for every $\varepsilon > 0$, there exists $0 < \eta(\varepsilon) < \varepsilon$ such that given sequences $(T_k) \subset S_{\mathcal{L}(X,Y)}$ and $(x_k) \subset S_X$, and a convex series $\sum_{k=1}^\infty \alpha_k$ such that
$$\left\| \sum_{k=1}^\infty \alpha_k T_k x_k \right\| > 1 - \eta(\varepsilon),$$
there exist a subset $A \subset \conjN$, $y^* \in S_{Y^*}$, and sequences $(S_k) \subset S_{\mathcal{L}(X,Y)}$, $(z_k) \subset S_X$ satisfying the following:
\begin{enumerate}
    \item[(1)] $\sum_{k \in A} \alpha_k > 1 - \varepsilon$,
    \item[(2)] $\|z_k - x_k\| < \varepsilon$ and $\|S_k - T_k\| < \varepsilon$ for all $k \in A$,
    \item[(3)] $y^*(S_k z_k) = 1$ for every $k \in A$.
\end{enumerate}
\end{dfn}

The authors \cite{KLM15} provided several examples of pairs $(X,Y)$ of Banach spaces that satisfy the generalized AHSP. For instance, if both $X$ and $Y$ are finite-dimensional, then the pair $(X,Y)$ has the generalized AHSP. Moreover, for any Hilbert space $H$, if the pair $(X,H)$ has the BPBp, then it also has the generalized AHSP. Later, in 2022, Kim and Lee \cite{KL22} studied this property for $Y$-valued function spaces over a base algebra $A$.

Kim et al. \cite{KLM15} also presented some important observations that will be useful in the present paper.

\begin{remark}\cite[Remark 5]{KLM15}
Let $X, Y$ be Banach spaces.
\begin{enumerate}
    \item [(a)] To show that the pair $(X,Y)$ has the generalized AHSP it is enough to check the conditions for every finite (but of arbitrarily length) convex series, with the same function $\varepsilon \mapsto \eta(\varepsilon)$ for all lengths.
    \item[(b)] In the definition of the generalized AHSP we may consider sequences $(T_k) \subset B_{\mathcal{L}(X,Y)}$ and $(x_k) \subset B_X$ (with a small change in the function $\eta(\varepsilon)$).
    \item[(c)] If $(X,Y)$ has the generalized AHSP, then $(X,Y)$ has the BPBp.
    \item[(d)] Also, if $(X,Y)$ has the generalized AHSP, then $Y$ has the AHSP. 
\end{enumerate}
\end{remark}

Given a set $\Gamma$ and a Banach space $X$, we write $B(\Gamma, X)$ to denote the Banach space of all bounded functions from $\Gamma$ into $X$, endowed with the supremum norm. We will consider several of its closed subspaces.

For a compact Hausdorff space $K$, we denote by $\mathcal{C}(K, X)$ the space of all continuous functions from $K$ into $X$. For a locally compact Hausdorff space $L$, we write $\mathcal{C}_0(L, X)$ for the Banach space of all continuous functions from $L$ into $X$ that vanish at infinity. For a completely regular space $\Omega$, we denote by $\mathcal{C}_b(\Omega, X)$ the Banach space of all bounded continuous functions from $\Omega$ into $X$.

Let $Z$ be any of the above subspaces of $B(\Gamma, X)$ and $t \in \Gamma$. We write $\delta_t(f) = f(t)$ for every $f \in Z$.

In this paper, we investigate pairs $(X,Y)$ that satisfy the BPBp and/or the generalized AHSP. For a locally compact Hausdorff space $L$, we show that the pair $(X,Y)$ has the BPBp whenever $(X, \mathcal{C}_0(L,Y))$ has the BPBp. However, the converse does not hold. We prove that if the pair $(X,Y)$ has the generalized AHSP and $\mathcal{L}(X,Z) = \mathcal{K}(X,Z)$, then the pair $(X,Z)$ has the generalized AHSP when $Z$ is one of the following spaces: $\mathcal{C}(K,Y)$ for any compact Hausdorff space $K$; $\mathcal{C}_0(L,Y)$ for any locally compact Hausdorff space $L$; and $\mathcal{C}_b(\Omega,Y)$ for any completely regular space $\Omega$.  We use techniques from \cite{ACKLM15} and \cite{CGKM14} to obtain these results.

\section{Results}

We begin this section by recalling a very interesting result from \cite{ACKLM15}.

\begin{theorem}\textnormal{\cite[Proposition 2.8]{ACKLM15}}\label{Prop2.8}
    Let $X$, $Y$ be Banach spaces and let $K$ be a compact Hausdorff space. If the pair $(X, \mathcal{C}(K,Y))$ has the BPBp, then the pair $(X,Y)$ has the BPBp.
\end{theorem}

The converse of the above theorem is false. Indeed, if $K = [0,1]$ and $Y=\conjR$ or $\conjC$, then $\mathcal{C}(K,Y)$ does not have the Lindenstrauss property B, as shown in \cite[Theorem A]{S83}, even though the pair $(X,Y)$ has the BPBp for every Banach space $X$, by the Bishop-Phelps theorem. However, the converse does hold for the generalized AHSP when $\mathcal{L}(X,\mathcal{C}(K,Y)) = \mathcal{K}(X,\mathcal{C}(K,Y))$. This result was demonstrated by Kim and Lee \cite{KL22} in the complex case. Next, we present an alternative proof valid for both the real and the complex case.

\begin{theorem}\label{Theo3.15}
    Let $X$, $Y$ be Banach spaces and let $K$ be a compact Hausdorff space. If the pair $(X,Y)$ has the generalized AHSP and $\mathcal{L}(X,\mathcal{C}(K,Y)) = \mathcal{K}(X,\mathcal{C}(K,Y))$, then the pair $(X, \mathcal{C}(K,Y))$ has the generalized AHSP. Equivalently, if the pair $(\ell_1(X),Y)$ has the BPBp and $\mathcal{L}(X,\mathcal{C}(K,Y)) = \mathcal{K}(X,\mathcal{C}(K,Y))$, then the pair $(\ell_1(X), \mathcal{C}(K,Y))$ has the BPBp.
\end{theorem}

\begin{proof}
    Since the pair $(X,Y)$ has the generalized AHSP, it follows that the pair $(\ell_1(X),Y)$ has the BPBp. As $\mathcal{L}(X,\mathcal{C}(K,Y)) = \mathcal{K}(X,\mathcal{C}(K,Y))$, we have $\mathcal{L}(X,Y) = \mathcal{K}(X,Y)$. Thus, by \cite[Lemma 3.10]{DGMM18}, the pair $(\ell_1(X),Y)$ has the BPBp for compact operators (see \cite[Definition 1.4]{DGMM18}). Applying \cite[Theorem 3.15]{DGMM18}, we obtain that the pair $(\ell_1(X), C(K, Y))$ has the BPBp for compact operators. Finally, using \cite[Lemma 3.10]{DGMM18} once more and the fact that $\mathcal{L}(X,\mathcal{C}(K,Y)) = \mathcal{K}(X,\mathcal{C}(K,Y))$, we conclude that the pair $(\ell_1(X), \mathcal{C}(K,Y))$ has the BPBp. Therefore, the pair $(X, \mathcal{C}(K,Y))$ has the generalized AHSP.
\end{proof}

Considering $\mathcal{C}_0(L,Y)$ instead of $\mathcal{C}(K,Y)$, we obtain analogous results.

\begin{theorem}\label{t1}
    Let $X$, $Y$ be Banach spaces and let $L$ be a locally compact Hausdorff space. If the pair $(X, \mathcal{C}_0(L,Y))$ has the BPBp, then the pair $(X,Y)$ has the BPBp.
\end{theorem}

\begin{proof}
    Suppose that $(X,\mathcal{C}_0(L,Y))$ has the BPBp with a function $\varepsilon \mapsto \eta(\varepsilon)$. Fix $\varepsilon>0$ and consider $T \in S_{\mathcal{L}(X,Y)}$ and $x_0 \in S_X$ such that
    $$\left\| T x_0 \right\| > 1-\eta\left(\frac{\varepsilon}{2}\right).$$

    Let $t_0 \in L$. Since $L$ is locally compact, there exist a compact set $K$ such that $t_0 \in \interior K \subset K$. Define $U = \interior K$. By hypothesis, $L$ is a completely regular space. Hence, there is a continuous function $\phi: L\rightarrow [0,1]$ such that $\phi(t_0)=1$ and $\phi(t)=0$ for every $t \in L\setminus U$.

    Consider the bounded linear operator $R: X \rightarrow \mathcal{C}_0(L, Y)$ defined by $Rx(t)=\phi(t)Tx$ for all $x\in X$ and $t\in L$. Thus,
    $$\|R\| = \sup_{x \in B_X} \sup_{t \in L} \left\| \phi(t)Tx \right\| = \sup_{t \in L} |\phi(t)| \cdot \sup_{x \in B_X} \| Tx \| = 1.$$
    Notice that 
    $$\left\| R x_0(t) \right\| = |\phi(t)| \| T x_0 \| \leq \| T x_0 \|.$$ Since $ \left\| R x_0(t_0) \right\| = \left\| T x_0 \right\|$, we have
    $$\left\| R x_0 \right\| = \left\| T x_0 \right\| > 
    1-\eta\left(\frac{\varepsilon}{2}\right).$$
    By the assumption, there exist $S \in S_{\mathcal{L}(X,\mathcal{C}_0(L,Y))}$ and $u_0 \in S_X$ satisfying
    $$\|Su_0\| = 1, \quad \|u_0 - x_0\| < \frac{\varepsilon}{2} \quad \textnormal{ and } \quad \|S - R\| < \frac{\varepsilon}{2}.$$
    In particular, there is $t_1\in L$ such that $\| S u_0(t_1) \| = \| Su_0 \| = 1$. 
    Now, define the bounded linear operator $\widetilde {S}: X \rightarrow Y$ by $\widetilde{S}=\delta_{t_1} \circ S$. Hence, 
    $$\|\widetilde{S}\| = \sup_{x \in B_X} \left\| Sx(t_1) \right\| =  \| S u_0(t_1) \| = \| \widetilde{S}u_0 \| = 1.$$
    Besides that,
    \begin{align*}
        1-\left\|R u_0(t_1)\right\| &= \|S u_0(t_1)\|-\|R u_0(t_1)\|\\
        &\leq \|S u_0(t_1)- R u_0(t_1)\|\\
        &\leq \left\| S-R\right\|<\frac{\varepsilon}{2}.
        \end{align*}
    This implies that $1-\frac{\varepsilon}{2}<\left\|Ru_0(t_1)\right\|$. That is, $\phi(t_1)Tu_0\neq 0$. The only possibility is that $t_1\in U$, and
    $$1-\frac{\varepsilon}{2} < \left\|Ru_0(t_1)\right\| = \left\|\phi(t_1)Tu_0\right\| \leq \phi(t_1).$$
    Finally, if $x\in B_X$ then
    \begin{align*}
    \left\|\widetilde{S}x-Tx\right\| &= \left\|(\delta_{t_1}\circ S)x-Tx\right\|\\
    &\leq \left\| Sx(t_1)-Rx(t_1)\right\|+\left\| Rx(t_1)-Tx\right\|\\
    &\leq \left\| S-R\right\|+\left\lvert \phi(t_1)-1\right\rvert\left\| Tx\right\|\\
    &\leq \left\| S-R\right\|+1-\phi(t_1)\\
    &<\left\| S-R\right\|+\frac{\varepsilon}{2}.
    \end{align*}
    Therefore, $\left\| \widetilde{S}-T\right\| \leq \left\| S-R\right\| + \frac{\varepsilon}{2}<\varepsilon$, which completes our proof.
\end{proof}

As in the case of Theorem \ref{Prop2.8}, the converse of the above result is false, by the same argument presented after Theorem \ref{Prop2.8}. Nevertheless, we obtain the following positive result.

\begin{theorem}\label{t2}
    Let $X$, $Y$ be Banach spaces and let $L$ be a locally compact Hausdorff space. If the pair $(X,Y)$ has the generalized AHSP and $\mathcal{L}(X,\mathcal{C}_0(L,Y))=\mathcal{K}(X,\mathcal{C}_0(L,Y))$, then the pair $(X, \mathcal{C}_0(L,Y))$ has the generalized AHSP. Equivalently, if the pair $(\ell_1(X),Y)$ has the BPBp and $\mathcal{L}(X,\mathcal{C}_0(L,Y))=\mathcal{K}(X,\mathcal{C}_0(L,Y))$, then the pair $(\ell_1(X), \mathcal{C}_0(L,Y))$ has the BPBp.
\end{theorem}

\begin{proof}
 Suppose that $(X,Y)$ has the generalized AHSP with a function $\varepsilon \mapsto \eta(\varepsilon)$. Fix $\varepsilon>0$ and consider sequences $(T_k)^{n}_{k=1}\subset B_{\mathcal{L}(X, \mathcal{C}_0(L,Y))}$, $(x_k)^{n}_{k=1}\subset B_{X}$ and a convex sum $\sum^{n}_{k=1}\alpha_k$ such that
    \begin{equation*}
        \left\| \sum_{k=1}^{n} \alpha_k T_k x_k \right\| >1-\eta\left(\frac{\varepsilon}{4}\right).
    \end{equation*}
Then, there is $t_0\in L$ such that
 \begin{equation}\label{i1}
        \left\| \sum_{k=1}^{n} \alpha_k T_k x_k(t_0) \right\|>1-\eta\left(\frac{\varepsilon}{4}\right).
    \end{equation}
    
For every $k\in \{1, \ldots, n \}$, we consider the bounded linear operator $R_k:X\rightarrow Y$ defined by $R_k=\delta_{t_0}\circ T_k$. Notice that $ \left\| R_k\right\|\leq 1$ for all $k\in \{1, \ldots, n \}$. By (\ref{i1}), 
    \begin{equation*}
         \left\| \sum_{k=1}^{n} \alpha_k R_k x_k \right\|=\left\| \sum_{k=1}^{n} \alpha_k  (\delta_{t_0}\circ T_k)x_k \right\|=\left\| \sum_{k=1}^{n} \alpha_k   T_kx_k(t_0) \right\|>1-\eta\left(\frac{\varepsilon}{4}\right).
    \end{equation*}
By the assumption, there exist a subset $A\subset \{1, \ldots, n \}$, $y^*\in S_{Y^*}$, and sequences $(S_k)^{n}_{k=1}\subset S_{\mathcal{L}(X,Y)}$ and $(z_k)^{n}_{k=1}\subset S_X$ satisfying the following:
    \begin{enumerate}
    \item[(1)] $\sum_{k \in A} \alpha_k > 1 - \frac{\varepsilon}{4}$,
    \item[(2)] $\|z_k - x_k\| < \frac{\varepsilon}{4}$ and $\|S_k - R_k\| < \frac{\varepsilon}{4}$ for all $k \in A$,
    \item[(3)] $y^*(S_k z_k) = 1$ for every $k \in A$.
    \end{enumerate}
We choose $\beta>0$ such that $\left\| R_k-S_k\right\|<\beta<\frac{\varepsilon}{4}$ for all $k\in A$, and consider $\delta=\frac{\varepsilon}{4}-\beta$. By the hypothesis, $\bigcup_{k\in A}\overline{T_k(B_X)}$ is a compact set. Then, for each $k\in A$, there exist $\widetilde{x}^k_1, \ldots, \widetilde{x}^k_{m_k} \in B_X$ such that $\bigcup_{k\in A}\overline{T_k(B_X)}\subset \bigcup_{k\in A}\bigcup^{m_k}_{j=1}B(T_k\widetilde{x}^k_j, \delta)$, where $B(T_k\widetilde{x}_j^k, \delta)$ denotes the open ball with center $T_k\widetilde{x}_j^k$ and radius $\delta$. We define the set
\begin{equation*}
    U=\bigcap_{k\in A}\bigcap^{m_k}_{j=1}\left\{t\in L:\, \left\| T_k\widetilde{x}^k_j(t)-S_k\widetilde{x}^k_j\right\|<\beta \right\}.    
\end{equation*}
As $U$ is a finite intersection of open subsets of $L$, then it is open. Furthermore, for every $k\in A$ and $j\in \{1,\ldots, m_k\}$, we see that 
$$\left\| T_k\widetilde{x}^k_j(t_0)-S_k\widetilde{x}^k_j\right\|=\left\| (\delta_{t_0}\circ T_k)\widetilde{x}^k_j-S_k\widetilde{x}^k_j\right\|=\left\| R_k\widetilde{x}^k_j-S_k\widetilde{x}^k_j\right\|\leq \left\| R_k-S_k\right\|<\beta.$$ Which implies that $t_0\in U$. Thus, there exists a compact set $K$ such that $t_0\in \interior K \subset K\subset U$. We consider $V=\interior K$. There is a continuous function $\phi: L \rightarrow [0,1]$ such that $\phi(t_0)=1$ and $\phi(t )=0$ if $t\in L\setminus V$.

For every $k\in A$, we define the bounded linear operator $\widetilde {S}_k: X\rightarrow \mathcal{C}_0(L,Y)$ by $$\widetilde {S}_kx=\phi S_kx+(1-\phi)T_kx.$$
We claim that $\left\| \widetilde {S}_k\right\|=1$ for all $k\in A$. Indeed, given $t\in L$ and $x\in B_X$, we have
 \begin{align*}
   \left\| \widetilde {S}_kx(t)\right\| &= \left\| \phi(t) S_kx+(1-\phi(t))T_kx(t)\right\|\\
   &\leq \phi(t)\left\| S_kx\right\|+(1-\phi(t))\left\| T_kx\right\|\\
   &\leq \phi(t)+1-\phi(t)=1
 \end{align*}
for all $k\in A$. We define the functional $\psi:\mathcal{C}_0(L,Y)\rightarrow \mathbb{K}$ by $\psi=y^*\circ \delta_{t_0}$. Notice that $\|\psi\| \leq 1$. 
Furthermore, 
\begin{align*}
\psi(\widetilde{S}_kz_k) &=y^*(\widetilde{S}_kz_k(t_0))\\
&=y^*(\phi(t_0)S_kz_k+(1-\phi(t_0))T_kz_k(t_0))\\
&=y^*(S_kz_k)=1
\end{align*}
for all $k\in A$. Therefore, $\left\| \psi\right\|=1$. We conclude that $\left\| \widetilde{S}_k\right\|=1$ for all $k\in A$ and also that the third condition of the Definition \ref{gahsp} is satisfied.

To finish the proof, let $x\in B_X$. Hence, for each $k\in A$, there is $j\in \{1, \ldots, m_k \}$ such that $\left\| T_kx-T_k\widetilde{x}^k_j\right\|<\delta$. If $t\in V$, then $\left\| T_k\widetilde{x}_j^k(t)-S_k\widetilde{x}_j^k\right\|<\beta$ for all $k\in A$ and for all $j\in \{1, \ldots, m_k \}$. This implies that

 \begin{align*}
     \left\| T_kx(t)-\widetilde{S}_kx(t)\right\| &= \left\| T_kx(t)-(\phi(t) S_kx+(1-\phi(t))T_kx(t))\right\|\\
     &= \left\| T_kx(t)-\phi(t) S_kx-T_kx(t)+\phi(t)T_kx(t)\right\|\\
     &= \left\| \phi(t)(T_kx(t)-S_kx)\right\|\\
     &\leq \left\| T_kx(t)-S_kx\right\|\\
     &\leq \left\| T_kx(t)-T_k\widetilde{x}^k_j(t)\right\|+\left\| T_k\widetilde{x}^k_j(t)-S_k\widetilde{x}^k_j\right\| +\left\| S_k\widetilde{x}^k_j-R_k\widetilde{x}^k_j\right\|\\
     &\hspace{0.4cm}+\left\| R_k\widetilde{x}^k_j-R_kx\right\|+\left\| R_kx-S_kx\right\|\\
     &< \left\| T_kx-T_k\widetilde{x}^k_j\right\|+\beta+ \left\| S_k-R_k\right\|+\left\| T_k\widetilde{x}^k_j(t_0)-T_kx(t_0)\right\|+\left\| S_k-R_k\right\|\\
     &< 2\delta+3\beta\\
     &= 2\left(\frac{\varepsilon}{4}-\beta \right)+3\beta=\frac{\varepsilon}{2}+\beta<\frac{3\varepsilon}{4}.
 \end{align*}
On the other hand, if $t\in L\setminus V$, then $\phi(t)=0$. In this case, $\left\| T_kx(t)-\widetilde{S}_kx(t)\right\|=0$. Thus, $\left\| T_k-\widetilde{S}_k\right\|\leq \frac{3\varepsilon}{4}<\varepsilon$ for all $k\in A$.
\end{proof}

An analogous result also holds for $\mathcal{C}_b(\Omega,Y)$ when $\Omega$ is a completely regular space. The proof is very similar and requires only minor modifications. The converse of this result remains an open question.
    
\begin{theorem}
    Let $X$, $Y$ be Banach spaces and let $\Omega$ be a completely regular space. If the pair $(X,Y)$ has the generalized AHSP and $\mathcal{L}(X,\mathcal{C}_b(\Omega,Y))=\mathcal{K}(X,\mathcal{C}_b(\Omega,Y))$, then the pair $(X, \mathcal{C}_b(\Omega,Y))$ has the generalized AHSP. Equivalently, if the pair $(\ell_1(X),Y)$ has the BPBp and $\mathcal{L}(X,\mathcal{C}_b(\Omega,Y))=\mathcal{K}(X,\mathcal{C}_b(\Omega,Y))$, then the pair $(\ell_1(X), \mathcal{C}_b(\Omega,Y))$ has the BPBp.
\end{theorem}

\

\noindent\textbf{Competing Interests} \ The authors report there are no competing interests to declare.

\

\noindent\textbf{Funding} \ The second author was partially supported by FAPEMIG RED-00133-21 and FAPEMIG APQ-01853-23.

\


\end{document}